if one side of a square is increased by 3 inches and the adjacent side is decreased by 2 inches, a rectangle is formed whose are is equal to the area of the square. Find a side of the square
Answer by ankor@dixie-net.com(22740)

if one side of a square is increased by 3 inches and the adjacent side is decreased by 2 inches, a rectangle is formed whose are is equal to the area of the square.
Find a side of the square
:
Let x = the side of the square
then
x^2 = the area of the square
:
(x+3) = length of the rectangle
and
(x-2) = width of the rectangle
then
(x+3)*(x-2) = x^2 + x - 6, the area of the rectangle
:
Rect area = square area
x^2 + x - 6 = x^2
x^2 - x^2 + x = 6
x = 6 inches is the side of the square
:
:
See if this flies:
(6+3)*(6-2) = 6^2
9 * 4 = 36